# This version: December 27, 2025)

### v2

@@ 1 Introduction
The main solution method for this problem is the first-order approach, which assumes that only local deviations in $a$ are binding, and yields a simple formula for the optimal contract. [^1] See [holmstrom1978incentives], the excellent survey [georgiadis2022contracting], and recent work in [conlon2009two, kadan2017existence, chade2020no, chaigneau2022should, castro2024disentangling]. Because of tractability, a vast literature simply assumes that the first-order approach is valid, and many applied papers make restrictive assumptions to avoid the issue of non-local deviations.222[mirrlees1999] showed that the first-order approach is not always valid. Examples of papers assuming the first-order approach include [jewitt2008moral, moroni2014existence, chaigneau2022should, castro2024disentangling]. Virtually all early work, including the seminal papers by [holmstrom1978incentives, holmstrom1979moral], and [zeckhauser1970medical] assumes it. [holmstrom1979moral] notes that “one has to assume that the agent’s optimal choice of action is unique for the optimal […] This assumption seems very difficult to validate […] and regrettably we have to leave the question about its validity open.” In applied work, restrictive assumptions such as linear contracts or binary effort are often made to avoid this issue. [edmans2009multiplicative] is an example of an important recent paper using both binary effort and linear contracts.
Unfortunately, existing sufficient conditions for the first-order approach to be valid are restrictive. The seminal papers are [rogerson1985] and [jewitt1988justifying], followed by an extensive literature. [^3] Important generalizations include [sinclair1994first, conlon2009two, jung2015information, chade2020no], and [jung2024proxy]. [chaigneau2022should, chaigneau2024theory] develop sufficient conditions with limited liability. [kadan2017existence] summarize the general view that “conditions facilitating the first-order approach are typically quite demanding.” The key issues are elegantly explained by [chaigneau2022should] and [conlon2009two], who says “Unfortunately, the Jewitt conditions are tied to the concavity, not only of the technology, but also of the payment schedule. Thus, there are interesting cases where the Jewitt conditions should fail, since the payment schedule is often not concave. For example, managers often receive stock options, face liquidity constraints.”444[conlon2009two] continues: “ I believe that it is natural to expect the first-order approach itself to fail in such cases, since the agent’s overall objective function will tend to be nonconcave […]. The fact that CDFC and CISP imply concavity of the agent’s payoff, regardless of the curvature of [the payment], then suggests that the CDFC/CISP conditions are very restrictive, not that the first-order approach is widely applicable.”
@@
Section 2 explains our results through simple examples and reconciles our findings with the literature. Section 3 gives definitions. Section 4 states Theorem 1. Section 5 outlines the proof. Section 6.1 provides a broad range of closed-form solutions for optimal contracts. Section 6.2 shows that optimal contracts are piecewise linear option contracts for log utility and output distributions in an exponential family with linear sufficient statistic (this has been independently discovered by [opp2025moral]). Section 6.3 provides an algorithm for finding the optimal contracts. Section 6.4 gives counter-examples and discusses limitations of our results. Section 6.5 discusses the relationship between our results and the literature.
@@ 2 Clarifying Example
The first-order approach fails with a reservation wage of $\mathdollar 1,000$, but is valid for all the illustrated positive reservation wages. This is an illustration of our main point, that the first-order approach is broadly valid in most interesting cases. Moreover, this is representative of a more general phenomenon, and not simply this example. Figures 2 and 3 consider a different distribution (the discrete Poisson distribution) and a different utility function (CARA). The same pattern holds, as it does in a multitude of other examples.
@@
Theorem 1 does not contradict existing results in the literature. The low reservation utility confirms the literature consensus that it is hard to guarantee that the first-order approach is valid is correct – if we insist that the first-order approach is valid for all reservation utilities. We see that, even in the Gaussian-log utility case, the first-order approach fails. Thus, any condition guaranteeing the first-order approach for all reservation utilities rules out basic examples and is therefore quite restrictive. The examples formally demonstrate the point made by [chaigneau2022should] and [conlon2009two], that requiring the first-order approach to be valid for all reservation utilities excludes many interesting cases.
As a corollary to Theorem 1, the problem is often tractable because we can use standard first-order approach arguments dating back to [holmstrom1978incentives]. We show that an optimal contract exists, is unique, and optimal wages have a simple formula:
@@ 5.1 Solution to the Relaxed Problem
We first show that the relaxed cost minimization problem has an almost everywhere unique solution, and that this solution has a simple formula in terms of Lagrange multipliers. This section is a minor extension of standard results on the relaxed problem going back to [holmstrom1978incentives] and [jewitt2008moral].
@@ 6.3 Numerical Methods
To see this, consider a relaxed cost minimization problem where we also include a finite number of global incentive compatibility constraints at a vector of actions $\boldsymbol{\hat{a}}=(\hat{a}_{1},\ldots,\hat{a}_{n})$. The Lagrangian from equation (2) becomes
@@
1
2
[…]
4
[…]
6
7repeat
@@
11
12 $\mathtt{best\_deviation},\;\mathtt{deviation\_gain}\leftarrow\textnormal{{FindBestDeviation(}}\textnormal{\emph{$v,\,a_{0}$}}\textnormal{{)}}$
13
14 if $\mathtt{deviation\_gain}>\mathtt{tolerance}$ then
@@
16
17   end if
18
19until $\mathtt{deviation\_gain}\leq\mathtt{tolerance}$
20return $v$
@@
Note that the dual maximization step is not a convex program, so this algorithm does not guarantee global optimality. In practice, it solves all the examples in the paper, and each run takes in the order of 1.5 to 8 milliseconds. Performance seems to be due to three factors. First, the algorithm often ends in the first iteration because of Theorem 1. Second, dual calls are fast because of the analytic formulas and caching. Third, dual maximization is fast because of low dimensionality, analytic gradients and warm starts.
@@ 6.4 Discussion of Assumptions and Counter-Examples
Limited liability. Limited liability is crucial to guarantee existence of an optimal contract. The intuition is clear from the Lagrangian’s first order condition in equation (FOC). When the right-hand side is negative, it is optimal to give the agent the lowest possible $v(y)$. This only has a solution with both limited liability and $u(0)>$-$\infty$. This is why with log utility $u(x)=\log(w_{0}+x)$ our assumptions require $w_{0}>0$. Under our assumption that the score is unbounded below, a negative right-hand side happens for any $\mu>0$. These existence issues are well-known since [mirrlees1999]. Limited liability is not important for guaranteeing the validity of the first-order approach, and in fact the kink in the optimal contract makes it more difficult to guarantee the validity of the first-order approach (see proof of proposition 2).
@@ 6.5 Discussion of the Literature
First Order Approach Literature. There are three generations of contributions to the first-order approach. The first generation was started by [mirrlees1999] (circulated in 1975), who showed that the first-order approach does not hold generally. Prior to [mirrlees1999], the first-order approach was simply assumed with little justification. Subsequent papers became aware that they had to assume validity of the first-order approach [holmstrom1978incentives], and developed results that do not depend on the first-order approach [grossman1983, araujo2001general].
The seminal papers in the second generation are [rogerson1985] and [jewitt1988justifying]. They provided sufficient conditions for the first-order approach. Rogerson’s conditions are well known to be strict, ruling out many natural distributions, including the Gaussian. Jewitt’s conditions are more general, and include many interesting cases. Jewitt’s main assumption is implicit, in that he requires his conditions to hold for all output levels, in particular ruling out limited liability. [chaigneau2022should] give a detailed explanation of how the Jewitt conditions are not applicable in the case of limited liability and survey the related literature. But limited liability is crucial to guarantee existence (see Section 6.4). In many examples that satisfy the Jewitt conditions, an optimal contract does not exist. This includes the Gaussian - log utility example without limited liability. At face, value, the Jewitt conditions are satisfied. But the relaxed cost minimization problem does not have a solution. Thus, the first-order approach is “valid”, since any of the non-existent solutions satisfies any property. But this validity is vacuous. This is similar to the colorful logical example that the statement “all unicorns are pink” is true, since unicorns do not exist. The second-generation papers have been influential, and much of the theoretical and applied work to this day assumes these conditions.
Our own results are inspired by [jewitt1988justifying], and we build on his insight. We impose similar conditions on the part of the contract where limited liability does not bind, while allowing for limited liability. The proof of Theorem 1 shows that, with high reservation utility, the limited liability constraint binds with low probability. We then show, in similar lines to [jewitt1988justifying], that the agent’s problem becomes concave.
There is a third generation of work extending Jewitt’s approach to multidimensional actions and richer information structures [conlon2009two, jung2015information, chade2020no, chaigneau2022should, jung2024proxy]. We believe our methods can be adapted to such environments, but we do not pursue that here.
@@
Linear Contracts and Option Contracts. A long literature seeks conditions under which optimal contracts are simple, such as linear contracts. The classic paper is [holmstrom1987aggregation], who rationalize linear contracts with CARA utility in a continuous time dynamic model with Brownian output process. [innes1990limited] rationalizes piecewise linear contracts with a risk neutral agent, limited liability, and a monotonicity restriction on the wage function. Notable contributions include [chassang2013calibrated] and [carroll2015robustness] who rationalize linear contracts with risk neutral agents and a robust maxmin notion of optimality.
@@
The most closely related work is the contemporaneous paper [opp2025moral]. He independently derived the linearity result from remark 1. As far as we know, [opp2025moral] is the first to derive piecewise linear contracts under a broad set of conditions in standard moral hazard with risk aversion. [opp2025moral] includes a complete analysis of the exponential distribution case. He characterizes optimal contracts with and without limited liability and when the optimal exists without limited liability. In particular, he shows that the optimal contract with limited liability is globally linear for low intended action but has a kink for high intended action. We have no overlap with these results.
[hemmer1999introducing] assumes a number of Jewitt’s conditions, and log utility with no limited liability. Their corollary 2 part 2 then shows that the optimal contract is globally linear. It is important to highlight the difference. They prove that contracts are globally linear, whereas we prove that contracts are piecewise linear. While their result is stronger, optimal contracts often do not exist in their setting (section 6.4). Thus, [opp2025moral] and our paper are the first to derive piecewise linear contracts under a broad set of conditions, but [hemmer1999introducing] were the first to notice linearity of $k\circ g$ to the best of our knowledge. [chaigneau2015changes, chaigneau2017prudence] also attribute this observation to [hemmer1999introducing].
@@
High Reservation Utility. In an influential paper, [chade2020highstakes] consider the principal-agent problem in the case where reservation utility converges to infinity. This paper is one of our inspirations in considering the case of high reservation utility, although it deals with completely different issues. They assume the first-order approach is valid (their assumption 4), and give no results on validity of the first-order approach. Their main result is that the wage cost to the principal is asymptotically the same as if effort was observable, so that agency costs are asymptotically small. This is a strong result that depends on detailed assumptions that rule out many of our examples. As they note, “key assumptions driving our results are on the agent’s utility for income” … “they are not without bite: log utility is excluded, and indeed the convergence results fail in this setting.”
@@
[castro2024disentangling] assumes that the principal’s expected wage cost is convex in the intended action, which they motivate with a result by [chade2020highstakes] that this is true with high stakes. But [castro2024disentangling] does not otherwise use high reservation utility. They assume that the first-order approach is valid (their assumption 1). Neither paper assumes limited liability, and [castro2024disentangling] assumes that utility in unbounded below but uses a bounded below utility in their numerical results. [chade2025jeroen] consider a moral hazard problem with a set actions in an interval plus one non-ordered action. They consider the relaxed problem, assuming that the first-order approach is valid among the actions in the interval. They consider high stakes using the results in [chade2020highstakes] to show that the optimal contract with several different utility functions converges to essentially the same contract.
@@
Numerical Analysis. There are four general approaches to numerically solve the moral hazard problem without resorting to the first-order approach.
@@
The first approach is to include additional no-jump constraints in the relaxed cost minimization problem. This was introduced in the 1970s by [mirrlees1999] and improved by [araujo2001general]. The main limitation is that this relies on ad hoc tricks to decide which constraints to add (see [ke2018general] p. 1426). Motivated by this, [ke2018general] introduce a sandwich method where the additional constraint is added adversarially. This method works well in examples where part of the problem happens to be analytically tractable.
@@
The two most popular approaches start by discretizing the set of actions, so that there is only a finite number of global incentive compatibility constraints (following [grossman1983]). The linear programming approach maximizes over a joint distribution of actions, output levels, and utility from wages, all of which have to be in a discrete grid [prescott1999primer, su2007computation]. While this has the advantage of making the principal’s problem a linear program, the main difficulty is the curse of dimensionality: the number of variables is the product of the sizes of the three grids. To address this, [su2007computation] introduced the mathematical program with equilibrium constraints approach. This approach maximizes over the contract and a distribution over actions. So the number of variables is the number of possible output values plus the number of possible actions. These approaches are often applied to relatively small problems. For example, [su2007computation]’s main numerical example has two possible output levels.
The ideal paper to contrast our algorithm with these existing methods is [armstrong2007stock]. They model executive compensation with log normally distributed output. This is exactly the same signal structure as our examples with the Gaussian distribution, but with the principal payoff being an exponential function of the action. They use mathematical programming with equilibrium constraints. Their figure 3 displays some example optimal contracts. The figure displays some of the key features of optimal contracts, such as a region of zero payment followed by an increasing region following the appropriate functional form. These are impressive results far ahead of their time. However, their algorithm does not match closely the optimal contract far away from the middle of the distribution of output. Moreover, our algorithm is faster. They report that each solution to the principal’s problem takes several hours. Our algorithm takes 50-150 milliseconds in the Gaussian case (but takes 300 milliseconds in the harder $t$-distribution case).
@@
The fourth approach in the literature is to approximate the agent’s utility as a function of actions with a ratio of polynomials [renner2015polynomial]. [renner2015polynomial] then show that the principal’s problem can be reduced to a nonlinear program. Crucially, their method can give global optimality guarantees. They illustrate their method solving the log normal example from [armstrong2007stock].
@@
In terms of algorithms, our approach is most related to the no-jump approach. We simply use a reasonable heuristic to add constraints, while taking advantage of Lagrange duality, parametric assumptions for fast dual calculations with analytic gradients, caching, and warm starts. Our algorithm has the disadvantage of not giving a global optimality guarantee, and is thus complementary to the LP and polynomial approaches. For example, one could use our algorithm to obtain a solution candidate and check it with the LP or polynomial approaches.

### v3

# This version: )

@@ 1 Introduction
The main solution method for this problem is the first-order approach, which assumes that only local deviations in $a$ are binding, and yields a simple formula for the optimal contract. [^1] See [20], the excellent survey [16], and recent work in [14, 27, 6, 8, 4]. Because of tractability, a vast literature simply assumes that the first-order approach is valid, and many applied papers make restrictive assumptions to avoid the issue of non-local deviations.222[29] showed that the first-order approach is not always valid. Examples of papers assuming the first-order approach include [23, 30, 8, 4]. Virtually all early work, including the seminal papers by [20, 21], and [38] assumes it. [21] notes that “one has to assume that the agent’s optimal choice of action is unique for the optimal […] This assumption seems very difficult to validate […] and regrettably we have to leave the question about its validity open.” In applied work, restrictive assumptions such as linear contracts or binary effort are often made to avoid this issue. [15] is an example of an important recent paper using both binary effort and linear contracts.
Unfortunately, existing sufficient conditions for the first-order approach to be valid are restrictive. The seminal papers are [35] and [24], followed by an extensive literature. [^3] Important generalizations include [36, 14, 25, 6], and [26]. [8, 9] develop sufficient conditions with limited liability. [27] summarize the general view that “conditions facilitating the first-order approach are typically quite demanding.” The key issues are elegantly explained by [8] and [14], who says “Unfortunately, the Jewitt conditions are tied to the concavity, not only of the technology, but also of the payment schedule. Thus, there are interesting cases where the Jewitt conditions should fail, since the payment schedule is often not concave. For example, managers often receive stock options, face liquidity constraints.”444[14] continues: “ I believe that it is natural to expect the first-order approach itself to fail in such cases, since the agent’s overall objective function will tend to be nonconcave […]. The fact that CDFC and CISP imply concavity of the agent’s payoff, regardless of the curvature of [the payment], then suggests that the CDFC/CISP conditions are very restrictive, not that the first-order approach is widely applicable.”
@@
Section 2 explains our results through simple examples and reconciles our findings with the literature. Section 3 gives definitions. Section 4 states Theorem 1. Section 5 outlines the proof. Section 6.1 provides a broad range of closed-form solutions for optimal contracts. Section 6.2 shows that optimal contracts are piecewise linear option contracts for log utility and output distributions in an exponential family with linear sufficient statistic (this has been independently discovered by [31]). Section 6.3 provides an algorithm for finding the optimal contracts. Section 6.4 gives counter-examples and discusses limitations of our results. Section 6.5 discusses the relationship between our results and the literature.
@@ 2 Clarifying Example
The first-order approach fails with a reservation wage of $\mathdollar 1,000$, but is valid for all the illustrated positive reservation wages. This is an illustration of our main point, that the first-order approach is broadly valid in most interesting cases. Moreover, this is representative of a more general phenomenon, and not simply this example. Figures 2 and 3 consider a different distribution (the discrete Poisson distribution) and a different utility function (CARA). The same pattern holds, as it does in a multitude of other examples. Readers can explore more examples in our interactive demo: https://eduardomazevedo.github.io/azevedo-wolff-2025-site/.
@@
Theorem 1 does not contradict existing results in the literature. The low reservation utility confirms the literature consensus that it is hard to guarantee that the first-order approach is valid is correct – if we insist that the first-order approach is valid for all reservation utilities. We see that, even in the Gaussian-log utility case, the first-order approach fails. Thus, any condition guaranteeing the first-order approach for all reservation utilities rules out basic examples and is therefore quite restrictive. The examples formally demonstrate the point made by [8] and [14], that requiring the first-order approach to be valid for all reservation utilities excludes many interesting cases.
As a corollary to Theorem 1, the problem is often tractable because we can use standard first-order approach arguments dating back to [20]. We show that an optimal contract exists, is unique, and optimal wages have a simple formula:
@@ 5.1 Solution to the Relaxed Problem
We first show that the relaxed cost minimization problem has an almost everywhere unique solution, and that this solution has a simple formula in terms of Lagrange multipliers. This section is a minor extension of standard results on the relaxed problem going back to [20] and [23].
@@ 6.3 Numerical Methods
Consider a relaxed cost minimization problem where we also include a finite number of global incentive compatibility constraints at a vector of actions $\boldsymbol{\hat{a}}=(\hat{a}_{1},\ldots,\hat{a}_{n})$. The Lagrangian from equation (2) becomes
@@
$\boldsymbol{\hat{a}}\leftarrow\emptyset$;  $\boldsymbol{\hat{\mu}}\leftarrow\emptyset$;  $\lambda\leftarrow\lambda_{\text{init}}$;  $\mu\leftarrow\mu_{\text{init}}$
$\mathbf{y}_{\mathrm{grid}}\leftarrow\textnormal{{InitializeGrid}}()$
repeat
@@
 $\mathtt{best\_deviation},\;\mathtt{deviation\_gain}\leftarrow\textnormal{{FindBestDeviation(}}\textnormal{\emph{$v,\,a_{0}$}}\textnormal{{)}}$
 if $\mathtt{deviation\_gain}>\mathtt{tolerance}$ then
@@
   end if
until $\mathtt{deviation\_gain}\leq\mathtt{tolerance}$
return $v$
@@
Algorithm 1 solves all the examples in the paper. In our benchmarks, each run takes in the order of 3 to 14 milliseconds (see supplementary material). Performance seems to be due to three factors. First, the algorithm often ends in the first iteration because of Theorem 1. Second, dual calls are fast because of the analytic formulas and caching. Third, dual maximization is fast because of low dimensionality, analytic gradients and warm starts.
Algorithm 1 has two main drawbacks. The dual is not convex, so that the optimization step can get stuck in a local maximum. And the heuristic to add constraints might also get stuck.
To overcome these drawbacks, we implement a straightforward convex optimization solver. We discretize both the outcome and action spaces, which yields a problem with a convex objective and linear constraints. We solve this discretized problem using CVXPY with the Clarabel solver. This approach is computationally more expensive—taking between 230 and 240 milliseconds in our benchmarks—because it solves for the contract $v$ at every point of the outcome grid. However, it provides a global optimality guarantee. In our implementation, Algorithm 1 uses this convex solver as a fallback.
The main drawback of the convex solver is that the numerical optimal contract $v(y)$ is numerically unstable at points where the density $f(y|a_{0})$ is small. This is illustrated in Figure 6. The figure shows that both solvers reach essentially the same objective value. In the middle of the distribution, they reach essentially the same optimal contract. However, in the tails of the distribution, the convex solver optimal contract is numerically unstable.
Figure 6: Comparison of Algorithm 1 and convex program solvers.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$. Both solvers reach essentially the same expected wage cost. The convex program solution is numerically unstable in the tails where the density $f(y|a_{0})$ is small. Parameters are as in Figure 1 with $\sigma=10$. The convex solver uses a grid of 201 points for the outcomes and 200 points for the actions.
@@ 6.4 Discussion of Assumptions and Counter-Examples
Limited liability. Limited liability is crucial to guarantee existence of an optimal contract. The intuition is clear from the Lagrangian’s first order condition in equation (FOC). When the right-hand side is negative, it is optimal to give the agent the lowest possible $v(y)$. This only has a solution with both limited liability and $u(0)>$-$\infty$. This is why with log utility $u(x)=\log(w_{0}+x)$ our assumptions require $w_{0}>0$. Under our assumption that the score is unbounded below, a negative right-hand side happens for any $\mu>0$. These existence issues are well-known since [29]. Limited liability is not important for guaranteeing the validity of the first-order approach, and in fact the kink in the optimal contract makes it more difficult to guarantee the validity of the first-order approach (see proof of proposition 2).
@@ 6.5 Discussion of the Literature
First Order Approach Literature. There are three generations of contributions to the first-order approach. The first generation was started by [29] (circulated in 1975), who showed that the first-order approach does not hold generally. Prior to [29], the first-order approach was simply assumed with little justification. Subsequent papers became aware that they had to assume validity of the first-order approach [20], and developed results that do not depend on the first-order approach [17, 1].
The seminal papers in the second generation are [35] and [24]. They provided sufficient conditions for the first-order approach. Rogerson’s conditions are well known to be strict, ruling out many natural distributions, including the Gaussian. Jewitt’s conditions are more general, and include many interesting cases. Jewitt’s main assumption is implicit, in that he requires his conditions to hold for all output levels, in particular ruling out limited liability. [8] give a detailed explanation of how the Jewitt conditions are not applicable in the case of limited liability and survey the related literature. But limited liability is crucial to guarantee existence (see Section 6.4). In many examples that satisfy the Jewitt conditions, an optimal contract does not exist. This includes the Gaussian - log utility example without limited liability. At face, value, the Jewitt conditions are satisfied. But the relaxed cost minimization problem does not have a solution. Thus, the first-order approach is “valid”, since any of the non-existent solutions satisfies any property. But this validity is vacuous. This is similar to the colorful logical example that the statement “all unicorns are pink” is true, since unicorns do not exist. The second-generation papers have been influential, and much of the theoretical and applied work to this day assumes these conditions.
Our own results are inspired by [24], and we build on his insight. We impose similar conditions on the part of the contract where limited liability does not bind, while allowing for limited liability. The proof of Theorem 1 shows that, with high reservation utility, the limited liability constraint binds with low probability. We then show, in similar lines to [24], that the agent’s problem becomes concave.
There is a third generation of work extending Jewitt’s approach to multidimensional actions and richer information structures [14, 25, 6, 8, 26]. We believe our methods can be adapted to such environments, but we do not pursue that here.
@@
Linear Contracts and Option Contracts. A long literature seeks conditions under which optimal contracts are simple, such as linear contracts. The classic paper is [19], who rationalize linear contracts with CARA utility in a continuous time dynamic model with Brownian output process. [22] rationalizes piecewise linear contracts with a risk neutral agent, limited liability, and a monotonicity restriction on the wage function. Notable contributions include [12] and [3] who rationalize linear contracts with risk neutral agents and a robust maxmin notion of optimality.
@@
The most closely related work is the contemporaneous paper [31]. He independently derived the linearity result from remark 1. [31] proves that the first-order approach is valid with an exponential distribution for some parameter values. As far as we know, [31] is the first to derive piecewise linear contracts under a broad set of conditions in standard moral hazard with risk aversion. [31] includes a complete analysis of the exponential distribution case. He characterizes optimal contracts with and without limited liability and when the optimal contract exists without limited liability. In particular, he shows that the optimal contract with limited liability is globally linear for low intended action but has a kink for high intended action. We have no overlap with these results.
In the accounting literature, we know of two papers that showed that (potentially piecewise) linear contracts are optimal under certain assumptions including the first order approach. The earliest is [18]. They assume a number of Jewitt’s conditions, and log utility with no limited liability. Their corollary 2 part 2 then shows that the optimal contract is globally linear. They prove that contracts are globally linear, whereas we prove that contracts are piecewise linear. While their result is stronger, optimal contracts often do not exist in their setting (section 6.4). The excellent graduate textbook [13] chapter 19 p. 154 also notes that piecewise linear contracts are optimal assuming that the first order approach is valid. Thus, they get to the generally correct form of piecewise linear contracts as opposed to the globally linear result. But they assume the first-order approach, and do not prove existence. Thus, [31] and our paper are the first to derive piecewise linear contracts under a broad set of conditions, rather than assuming the first-order approach. But [18] were the first to notice linearity of $k\circ g$ to the best of our knowledge. [11, 10] also attribute this observation to [18].
@@
High Reservation Utility. In an influential paper, [5] consider the principal-agent problem in the case where reservation utility converges to infinity. This paper is one of our inspirations in considering the case of high reservation utility, although it deals with completely different issues. They assume the first-order approach is valid (their assumption 4), and give no results on validity of the first-order approach. Their main result is that the wage cost to the principal is asymptotically the same as if effort was observable, so that agency costs are asymptotically small. This is a strong result that depends on detailed assumptions that rule out many of our examples. As they note, “key assumptions driving our results are on the agent’s utility for income” … “they are not without bite: log utility is excluded, and indeed the convergence results fail in this setting.”
@@
[4] assumes that the principal’s expected wage cost is convex in the intended action, which they motivate with a result by [5] that this is true with high stakes. But [4] does not otherwise use high reservation utility. They assume that the first-order approach is valid (their assumption 1). Neither paper assumes limited liability, and [4] assumes that utility in unbounded below but uses a bounded below utility in their numerical results. [7] consider a moral hazard problem with a set actions in an interval plus one non-ordered action. They consider the relaxed problem, assuming that the first-order approach is valid among the actions in the interval. They consider high stakes using the results in [5] to show that the optimal contract with several different utility functions converges to essentially the same contract.
@@
Numerical Analysis. We now review existing approaches to numerically solve the moral hazard problem without the first-order approach. The review suggests that algorithm 1 offers a significant improvement. We are, to the best of our knowledge, the first to numerically solve the moral hazard problem for various realistic distributions without the first-order approach.
@@
The literature proposes four general approaches to numerically solve the moral hazard problem without the first-order approach. The first approach is to include additional no-jump constraints in the relaxed cost minimization problem. This was introduced in the 1970s by [29] and improved by [1]. The main limitation is that there is no general method to decide which constraints to add (see [28] p. 1426). Motivated by this, [28] introduce a sandwich method where the additional constraint is added adversarially. This method works well in some special cases that are analytically tractable, such as relative risk aversion with coefficient $1/2$ where the link function is piecewise linear.
@@
The two most popular approaches start by discretizing the set of actions, so that there is only a finite number of global incentive compatibility constraints (following [17]). The linear programming approach maximizes over a joint distribution of actions, output levels, and utility from wages, all of which have to be in a discrete grid [32, 33, 37]. While this has the advantage of making the principal’s problem a linear program, the main difficulty is the curse of dimensionality: the number of variables is the product of the sizes of the three grids. To address this, [37] introduced the mathematical program with equilibrium constraints approach. This approach maximizes over the contract and a distribution over actions. So the number of variables is the number of possible output values plus the number of possible actions.
We know of two papers who apply these approaches to realistic distributions. [33] considers a gaussian distribution with an improvement on the linear programming approach by using the Dantzig-Wolfe decomposition algorithm. [2] considers a log normal distribution, to study executive compensation, and use a mathematical program with equilibrium constraints approach. Both papers are far ahead of their time, as their numerical solutions display the key features of optimal contracts. They also report the wall time needed to solve the principal’s problem. [33] reports 51 minutes with 100 actions and 50 output levels. [2] reports that each solution takes "several hours" with 101 actions and 501 output levels. In our supplementary material benchmark table, algorithm 1 takes 93 to 356 milliseconds to solve similar Gaussian problems (without the need to discretize either actions or output levels – except for taking integrals). Naturally, these are not apples-to-apples comparisons due to improvements in hardware and implementation details. But the large performance gap suggests that algorithm 1 is substantially faster. This is consistent with algorithm 1 being significantly faster than the convex solver (see supplementary material). This is intuitive because algorithm 1 uses the analytic formulas from tables 1 and 2 and solves lower-dimensional problems.
@@
The fourth approach in the literature is to approximate the agent’s utility as a function of actions with a ratio of polynomials [34]. [34] then use results from polynomial optimization to show that the principal’s problem can be reduced to a nonlinear program. Crucially, their method can give global optimality guarantees. They solve the log normal example from [2] in the case of a restricted contract space, with a fixed payment, equity, and options. They generously shared their code. Our benchmarks suggest that their suggested nonlinear program takes about 100 seconds in Scipy’s SLSQP solver.
@@
Our approach is influenced by and builds upon this literature. Algorithm 1 is most related to the no-jump approach. We simply use a reasonable heuristic to iteratively add constraints, while taking advantage of Lagrange duality, parametric assumptions for fast dual calculations, analytic gradients, caching, and warm starts. Algorithm 1 has the disadvantage of not giving a global optimality guarantee, and is thus complementary to the LP and polynomial approaches. For example, one could use our algorithm to obtain a solution candidate and check it with the LP or polynomial approaches. Our convex programming approach is a natural modern extension of the linear programming and mathematical programming with equilibrium constraints approaches. Since the 2010s, there has been increasing availability of efficient convex solvers that are simple for non-experts to use, and it has become the standard tool for solving these types of problems in engineering. Our convex programming approach simply applies these new tools to the standard discretization approach from the literature going back to [17].